OpenStudy (anonymous):

Your computer-supply store sells two types of inkjet printers. The first, type A, costs $237 and you make a $22 profit on each one. The second, type B, costs $122 and you make a $19 profit on each one. You can order no more than 120 printers this month, and you need to make at least $2,400 profit on them. If you must order at least one of each type of printer, how many of each type of printer should you order if you want to minimize your cost? 69 of type A : 51 of type B 40 of type A : 80 of type B 51 of type A : 69 of type B 80 of type A : 40 of type B

hero (hero):

I could have sworn I helped you with this yesterday.

hero (hero):

You can't say this one is different.

OpenStudy (anonymous):

I don't think you did? I don't remember having a problem like this ?

hero (hero):

Hang on a minute.

OpenStudy (anonymous):

ok.

hero (hero):

I don't know how you managed to delete those questions you posted yesterday.

OpenStudy (anonymous):

I didn't , I don't even know how to delete them ?

hero (hero):

If we let: a = number of type A computers b = number of type B computers Then we can produce the following equations: Amount of Computers a + b ≤ 120 Cost of Computers: C(a,b) = 237a + 122b Profit from computers: 22a + 19b ≥ 2400 So now we take the amount equation and the profit equation and solve for a and b: a + b ≤ 120 22a + 19b ≥ 2400 Multiply the first equation by 22 to get: 22a + 22b ≤ 2640 22a + 19b ≥ 2400 Isolate 22a in both equations: 22a ≤ 2640 - 22b 22a ≥ 2400 - 19b Now set: 22a ≤ 22a 2400 - 19b ≤ 2640 - 22b Place like terms on the same side: 22b - 19b ≤ 2640 - 2400 3b ≤ 240 b ≤ 240/3 b ≤ 80 Thus a ≤ 40 since we know that a + b ≤ 120

hero (hero):

That's the same solution I typed up yesterday.

hero (hero):

That's why I was able to post it so quickly

hero (hero):

Thus that proves the question was posted yesterday

hero (hero):

Furthermore, since a = 40 and b = 80, we can find the total cost: C(a,b) = 237a + 122b C(40,80) = 237(40) + 122(80) = 19,240

OpenStudy (anonymous):

I do remember that . But, I swear I didn't delete anything ? and so you mean to say my teacher gave me the same question for yesterday & today ? lol

hero (hero):

After further review, it appears that someone else must have posted the same question.

hero (hero):

An @MeganDarling

OpenStudy (anonymous):

oh , but is it even possible to delete questions you have asked ?

hero (hero):

No, not anymore

hero (hero):

However, it is possible to alter your question.

OpenStudy (anonymous):

Oh , well thank you for helping me again .

OpenStudy (anonymous):

hi? haha

OpenStudy (anonymous):

hey , lol . are you in connections acad. ?